# Social learning moderates the tradeoffs between efficiency, stability, and equity in group foraging

### v1

@@
Ze-Xu Lia, M. Amin Rahimianb,∗ and Lei Fanga,c,d∗
@@
Social learning shapes collective search by influencing how individuals use peer information. Empirical and computational studies show that optimal information sharing, neither too localized nor too diffuse, can improve resource detection and group coordination. Building on these insights, we develop a randomized search model integrating social learning with area-restricted search (ARS) to examine how communication distance affects collective foraging outcomes. Our model incorporates three behavioral modes: exploration, exploitation, and targeted walk, to evaluate how groups balance local exploitation with global exploration. These modes are controlled by one key parameter $\rho$ that balances exploration and exploitation between the collective. We examine how communication distance affects group efficiency ($\eta$), its temporal variability ($\sigma$), and equity in resource distribution ($\epsilon$). We identified an intriguing trade-off between the three outcomes. As $\rho\to 0$, the agents explore the search field independently, leading to maximum collective exploration. With increasing $\rho$, agents prioritize exploiting patches that other members of the group identify before engaging in their own explorations. The $\eta$ is maximized while $\sigma$ is minimized at interior values of $\rho$ that effectively balance exploration and exploitation. As $\rho$ increases to the maximum range in the search field, the equality among individuals is maximized, but this comes at the cost of lower $\eta$ and $\sigma$ values. In addition, we introduce negative rewards to assess how social learning aids in risk mitigation and observe that search efficiency is optimized at lower values of $\rho$ with increasing number of negative targets or magnitude of negative rewards. Surprisingly, we can also use “fake” negative rewards to improve the efficiency of collective search by focusing the collective exploration effort away from empty regions of the search field. Our results clarify the trade-offs that shape adaptive collective behavior in complex environments.
Search is a critical behavior for mobile foragers, such as albatrosses [1], marine predators [2, 3], and hunter-gatherers [4, 5, 6], who must secure random and often sparsely distributed food resources in complex environments. Active search mode demands substantial allocation of limited energy budget and increases exposure to predators and environmental hazards. These challenges underscore the need for efficient search strategies that balance energy expenditure with resource acquisition.
A widely studied strategy for efficient search is the Lévy walk [1, 2, 3], a movement pattern that is closely related to but physically distinct from the mathematical construct of the Lévy flight. Both describe step lengths drawn from a power law distribution $P(l)\sim l^{-\mu}$ ($1<\mu\leq 3$), which produces clusters of short displacements interspersed with occasional very long steps, forming fractal patterns without a characteristic scale. In the idealized Lévy flight, the steps are instantaneous, allowing arbitrarily long jumps in zero time, whereas a Lévy walk couples each step length to the time required to traverse it at finite speed, making it the appropriate model for real organisms. Early empirical work often used the term Lévy flight when the measured trajectories were, in fact, finite-speed Lévy walk. In their seminal work [1], Viswanathan et al. (1996) test the Lévy foraging hypothesis with flight patterns of the wandering albatross. They posit that Lévy foraging represents an evolved, optimal strategy for searching in uncertain environments, a claim bolstered by Viswanathan et al. (1999), who mathematically demonstrate that power-law step length statistics can outperform other movement patterns, such as Brownian motion, when targets are sparsely distributed [7]. However, early enthusiasm was tempered by critiques, notably Edwards et al. [8], who expose statistical missteps in attributing the power-law scaling to some species, prompting the need for more rigorous methodologies. With improved analysis, the Lévy walk-like search has been confirmed in various systems, including T cells in the brain [9], swarming bacteria [10, 11], honeybees [12], aquatic marine predators [2, 3], and even hunter-gatherers [4].
Research on the Lévy foraging hypothesis suggests that animals searching for sparse and randomly distributed resources may benefit from movement patterns that approximate Lévy walks [7, 13, 14]. More recent work on intermittent, two-stage search shows that real foragers rarely follow a purely memoryless Lévy process. Food encounters often trigger area-restricted search (ARS) behavior, in which individuals switch from rapid, wide-ranging motion to slower, more tortuous paths near a resource patch [15, 16, 17, 18]. Modeling studies demonstrate that such encounter-driven ARS can outperform idealized Lévy walks by reducing the overall search time [16], increasing the encounter rates, and lowering the risk of starvation [15], especially when resources are scattered in patches.
In nature, many organisms search for food items collectively for survival. When an animal is not successful in locating food sources, it can instead locate other distant group members who are successful and join them. Collective search is a ubiquitous phenomenon observed across different species [19, 20, 21, 22], which harnesses the power of group dynamics to improve efficiency within environments where resources are often scarce, highly clustered, or ephemeral [23].
The efficiency of collective problem solving, however, depends critically on how information flows through social networks. Lazer & Friedman [24] demonstrated that network topology regulates the balance between exploration and exploitation; Mason & Watts [25] provided experimental evidence that efficient communication networks enhance group learning; and Becker et al. [26] revealed that patterns of social influence alter collective accuracy in “wisdom-of-crowds” tasks. Within this framework, our communication-range parameter $\rho$ serves as a spatial analogue of network connectivity, enabling direct examination of how information-exchange density regulates group performance. By linking ecological models of foraging with network-theoretic approaches to collective intelligence, we bridge behavioral ecology and social-network science, providing a mechanistic foundation for how connectivity mediates exploration–exploitation trade-offs.
At the heart of these dynamics lies social learning-the process by which individuals acquire and use information from others to guide behavior [27, 28]. The role of social learning in the collective search is multifaceted as it mediates how individuals process and act on information from their peers, profoundly affecting both individual and group outcomes. Empirical work on Mongolian gazelles, for instance, shows that vocal communication over intermediate distances optimizes foraging efficiency by guiding group members toward rich resource patches while avoiding the pitfalls of overly localized or excessively diffuse information sharing [29]. Similarly, computational models reveal that probabilistic interactions, in which foragers selectively move toward sites of successful peers, can generate scale-free movement patterns similar to Lévy walks, enhancing search efficiency in heterogeneous landscapes [27]. Two trade-offs have been found, exploration versus exploitation, and individual search versus social learning, to affect group efficiency [30]. Evolutionary perspectives further highlight that social learning can drive the coexistence of explorative and exploitative strategies within groups. However, excessive copying of others leads to suboptimal group efficiency unless it is mitigated by adaptive heuristics, such as area-restricted search [31].
Here, we use a deterministic (range-based) social learning model integrated with area-restricted search (ARS) to investigate the optimal communication distance at which groups can balance local exploitation and global exploration to maximize foraging efficiency across diverse target environments. In our model, we define three behavioral strategies: exploration, exploitation, and targeted walk to examine how the communication distance influences the efficiency of collective search. In addition, we investigate its impact on temporal patterns of efficiency and equitable resource distribution between individuals. Furthermore, we incorporated negative targets into our model to elucidate the effects of social learning on risk mitigation in foraging environments.
In this paper, we model group foraging by $N_{A}$ agents that start at random locations within the domain and follow a Lévy walk with $\mu=1.1$, executing an explorative strategy with longer steps; see Materials and Methods for details of the Lévy walk model with two modes of exploration and exploitation. Upon detecting a target, an agent will collect it and send a social signal sharing the location of the detected target with all other agents within a radius of $\rho$, which is the social learning parameter that determines the range of response to a social signal. Simultaneously, the agent that has just collected a target adjusts its search strategy from explorative to exploitative, changing $\mu$ from $1.1$ to $3$ to perform an area-restricted search. When an agent does not find any new targets within radius $R$ of its last collected target, it ends its area-restricted search and reverts to the explorative search strategy, changing its $\mu$ from $3$ to $1.1$ (Fig. 1).
Agents who have not received social signals (out of the social learning range $\rho$) or are conducting an area-restricted search continue to perform their own search strategy independently. In contrast, agents within the social learning range performing an explorative search ($\mu=1.1$) will stop their random walk and move towards the location where the social signal was sent (targeted walk). However, agents who are already doing a targeted walk will compare the distances between the two different destinations and choose the closest one. During the targeted walk, if agents encounter a new target, they will terminate the targeted walk and initiate an area-restricted search in its vicinity.
@@
Figure 1: A schematic of the collective search model. Agents begin by performing independent Lévy walks with $\mu=1.1$ (exploration, shown in purple) until they detect a target. Upon detection, the agent ceases social learning and switches to a Lévy walk with $\mu=3$ (exploitation, shown in blue) within a circular region of radius $R$ centered at the detected target, which is updated over time. Meanwhile, a social signal emitted by this agent attracts other agents within a region of radius $\rho$ to perform targeted walk towards the detected target (shown in orange). Agents outside this region do not receive the social signal and maintain their search mode.
@@ Results
We simulate our group foraging model by varying the social learning range from $\rho=0$ to $\rho=0.7$, which is the largest social learning range in our domain, using a periodic boundary condition. Considering that the targets are limited and nonregenerative, search efficiency will definitely drop as more targets are depleted. Each simulation will end when $30\%$ of the targets have been collected [30, 31]. We also vary the number of patches, keeping the total number of targets across all patches fixed at $6000$. Furthermore, we introduce negative targets, conceptualized as risks encountered by agents during the search, that negatively impact search efficiency. When an agent encounters a negative target, it can also send a social signal to inform agents within a radius $\rho$ of the presence of the negative target. Upon approaching known negative targets, agents exhibit active avoidance behaviors.
@@
The group efficiency, defined as $\eta=N_{collected}/N_{positive}T$, quantifies the rate at which agents collect positive targets. Here, $N_{collected}$ and $N_{positive}$ are the number of targets collected by agents and the total number of targets within the domain, respectively. $T$ is the total search time. As the social learning range ($\rho$) increases, it first increases and then decreases, with an optimal $\rho$ that yields the highest efficiency. (Fig. 2A). This phenomenon is robust across different numbers of patches, whether we fix the total number of targets or the number of targets per patch (see Supplementary Fig. 4). It can also be observed for target distributions with $\beta=1.1$ (see Supplementary Fig. 2) and $\beta=3$ (see Supplementary Fig. 3), as well as different group sizes when $N=10$ (see Supplementary Fig. 5). It is not surprising because Garg et al. [30] have revealed two trade-offs: the trade-off between exploitation and exploration in individual search controlled by the power-law exponent $\mu$, and the trade-off between individual search and social learning modulated by selectivity parameter. However, our clearer definition of exploration and exploitation can better explain the occurrence of the optimal $\rho$.
@@
Figure 3: Heterogeneity across time and space. (A) Time variation of instantaneous efficiency $\eta_{i}$ with $\rho=0.01$, $\rho=0.15$, and $\rho=0.7$ (showing one case). $\eta_{i}$ represents the number of collected targets per time step. (B) Probability density function (PDF) of $\eta_{i}$ indicating that low $\rho\leq 0.2$ limits $\eta_{i}$ due to less social learning, while higher $\rho$ ($0.2<\rho<0.5$) allows $\eta_{i}$ to reach greater values with an exponential distribution. For $\rho\geq 0.5$, the proportion of medium $\eta_{i}$ values decreases, though $\eta_{i}$ can achieve higher values with a heavy-tail distribution. (C) PDF of inter-event time, defining an event as when the instantaneous efficiency $\eta_{i}\geq 1$. The inset figure illustrates the variation of the burstiness parameter $B$ with $\rho$. The optimal $\rho$ corresponds to the lowest B, indicating a steady intake of resources and, consequently, a minimal starvation time. (D) PDF of number of collected targets by each agent, with an inset indicating that the standard deviation (representing equity) decreases as $\rho$ increases. All data is based on a scenario with 50 agents and $N_{+}=10$.
Agents in our model exhibit three modes: Levy work with $\mu=1.1$, Levy work with $\mu=3$, and targeted walks triggered by social learning. We interpret $\mu=1.1$ mode as global exploration, $\mu=3$ mode as area restricted exploitation, and the targeted walk as the operational cost associated with exploitation. Agents can adaptively switch between the three modes as needed. We measure the time spent on each mode, as shown in Fig. 2B. When $\rho=0$ (no social learning), targeted walks do not occur, and agents search independently, spending approximately 90% of their time on exploration. After encountering a target, an agent shifts from exploration to exploitation, which contributes to the remaining 10% of time spent on exploitation. As $\rho$ increases, targeted walks become more prevalent, triggering greater exploitation. Consequently, the time ratio of exploration decreases as $\rho$ rises. At the optimal $\rho$, the time ratio of exploitation reaches its maximum. This reflects an elegant balance between exploration and exploitation: enough exploitation sustains high current efficiency, while just enough exploration continually discovers new patches, preserving future opportunities and, in turn, overall efficiency. As $\rho$ increases further, targeted walks become more prevalent. However, the time ratio of exploitation decreases, indicating that an excessively high proportion of targeted walks leads to redundant search effort and reduces the exploitation ratio. This occurs because the number of targets is limited in each patch. When social signals attract most of the agents to a detected target, early-arriving agents can still collect targets, whereas late-arriving agents who travel for extended periods can only collect a few targets. Additionally, the remaining targets may be insufficient to support sustainable exploitation, causing late-arriving agents to revert to exploration soon. In other words, the social signal breaks the exploration of those late-arriving agents (far from the social signal) and lets them restart exploration from a region where most of the targets are consumed. Thus, an excessively large $\rho$, causes a high proportion of targeted walks, raises the cost and reduces the benefits of exploitation, disrupts the exploration process, and forces agents to explore regions where most targets are already consumed, thereby diminishing efficiency.
@@
To investigate the role of social learning and the resulting targeted walk, we construct a phase diagram in 2C, plotting the counts of targeted walks normalized to the total time in each case, against the averaged length of targeted walks. In this diagram, a higher count of targeted walks signifies increased interaction with social signals, which promotes exploitation and enhances efficiency. Conversely, a larger average length of targeted walks represents high cost, as it is the distance agents must travel before they can begin detecting a resource, which reduces overall efficiency. The slope of the curves in this phase diagram, therefore, represents the ratio of benefits to costs, specifically the gain from increased exploitation per unit of cost. Based on this, we hypothesize that the point where the slope of each curve reaches its maximum corresponds to the most efficient scenario. At this point, the benefit gained from targeted walks increases most rapidly relative to their cost. To verify this hypothesis, we compute the derivative of the curves from the phase diagram as shown in Fig. 2D, which reveals the slope. From the figure, a maximum value of the derivative represents the most efficient scenario. This derivative is smoothed using a Gaussian filter with $\sigma=1$ to reduce the noise.
@@
To highlight the burstiness of $\eta_{i}$, we plot its probability density functions (PDFs), as shown in Fig. 3B. The PDF reveals that a larger $\rho$ increases the proportion of higher $\eta_{i}$ values, resulting from social learning and agent aggregation. To further reveal the bursty pattern existing in the timeseries of instantaneous efficiency, we use the burstiness parameter introduced by Goh and Barabási [32]. Here, we defined an event as occurring when $\eta_{i}$ exceeds 0, while the interevent time $\tau$ represents the starvation time during which agents do not collect any targets. The burstiness parameter is quantified by $B\equiv(\sigma_{\tau}-m_{\tau})/(\sigma_{\tau}+m_{\tau})$, where $m_{\tau}$ and $\tau_{m}$ are the mean and standard deviation of $\tau$, respectively. Fig. 3C shows the PDFs of interevent times, where cases with high and low $\rho$ exhibit a larger proportion of starvation time, reflecting greater burstiness and leading to lower efficiency. The inset in Fig. 3C plots the burstiness parameter $B$ as a function of $\rho$, where $\rho$ values with lower $B$ correspond to the optimal $\rho$. Thus, the analysis of the time series and the derived burstiness provides another perspective on the optimal $\rho$, which not only exhibits the highest efficiency but also steady intake of resources and reduced burstiness in starvation time.

### v2

@@
Zexu Lia, M. Amin Rahimianb,∗ and Lei Fanga,c,d∗
@@
Collective foragers—from animals to robotic swarms—must balance exploration and exploitation to locate sparse resources efficiently. While social learning is known to facilitate this balance, how the range of information sharing shapes group-level outcomes remains unclear. Here, we develop a minimal collective foraging model in which individuals combine independent exploration, local exploitation, and socially guided movement. We show that foraging efficiency is maximized at an intermediate social learning range, where groups exploit discovered resources without suppressing independent discovery. This optimal regime also minimizes temporal burstiness in resource intake, reducing starvation risk. Increasing social learning range further improves equity among individuals but degrades efficiency through redundant exploitation. Introducing risky (negative) targets shifts the optimal range upward; in contrast, when penalties are ignored, randomly distributed negative cues can further enhance efficiency by constraining unproductive exploration. Together, these results reveal how local information rules regulate a fundamental trade-off between efficiency, stability, and equity, providing design principles for biological foraging systems and engineered collectives.
Efficient resource acquisition is a critical challenge for mobile foragers—including albatrosses [viswanathan1996levy], marine predators [sims2008scaling, sims2012levy], and human hunter-gatherers [raichlen2014evidence, venkataraman2017hunter, pacheco2019nahua]—who must locate spatially sparse and unpredictable food while expending limited energy and avoiding hazards. These constraints motivate strategies that balance wide-ranging exploration with localized exploitation.
A foundational framework for describing animal movement is the Lèvy walk [viswanathan1996levy, sims2008scaling, sims2012levy], in which step lengths follow a power-law distribution $P(l)\sim l^{-\mu}$ ($1<\mu\leq 3$), generating clusters of short displacements punctuated by rare long relocations. Although sometimes conflated with Lèvy flights—whose instantaneous steps lack physical realism—Lèvy walks capture finite-speed motion and have been supported empirically across diverse systems, from wandering albatrosses [viswanathan1996levy] to T cells [harris2012generalized], swarming bacteria [ariel2015swarming, ariel2017chaotic], honeybees [reynolds2007displaced], marine predators [sims2008scaling, sims2012levy], and human foragers [raichlen2014evidence]. Early claims of Lèvy scaling were tempered by statistical reanalyses [edwards2007revisiting], but subsequent improved methodology reaffirmed its relevance. While Lèvy strategies enhance encounters with sparse targets [viswanathan1999optimizing, viswanathan2008levy, humphries2014optimal], real foragers rarely follow memoryless step-length statistics. Upon finding prey, individuals typically switch to area-restricted search (ARS) characterized by slow, tortuous movements near resource patches [ross2018general, benichou2006two, benhamou2015ultimate, pyke2015understanding]. ARS reduces search time [benichou2006two], increases encounter rates, and lowers starvation risk [ross2018general], especially when resources are patchy.
However, many organisms do not forage alone. When unsuccessful individuals encounter other individuals exploiting rich patches, they can abandon independent search and join them. Such collective search is widespread across taxa [barnard1981producers, detrain2008collective, ginelli2015intermittent, gomez2022intermittent] and is especially advantageous when resources are scarce, clustered, or ephemeral [giraldeau2018social]. The efficiency of collective problem solving depends critically on how information flows through social networks. Network-theoretic analyses show that topology regulates the exploration–exploitation balance [lazer2007network]; experiments demonstrate that communication networks can enhance group learning [mason2012collaborative]; and social-influence structure shapes collective accuracy [becker2017network]. In this context, the communication radius $\rho$ can be viewed as a spatial analogue of network connectivity: small $\rho$ encourages independent search, while large $\rho$ increases conformity to shared discoveries. By linking ecological foraging models to social-network science, we provide a mechanistic perspective on how connectivity regulates group search.
At the core of these processes is social learning—the acquisition and use of information from others [bhattacharya2014collective, heins2024collective]. Social learning influences resource acquisition by tuning how individuals respond to peer discoveries. Empirical work on Mongolian gazelles shows that intermediate-distance vocal communication maximizes patch detection while avoiding over-aggregation [martinez2013optimizing]. Computational studies demonstrate that probabilistic copying can generate Lèvy-like statistics [bhattacharya2014collective], while theory reveals two fundamental trade-offs—exploration vs. exploitation and individual vs. social search—that govern efficiency [garg2022individual]. Evolutionary models further show that social learning promotes coexistence of diverse strategies, although excessive copying can harm group performance unless moderated by heuristics such as ARS [garg2024evolution]. Understanding these mechanisms is relevant not only for animal foraging but also for engineered multi-agent systems—such as swarm robotics and distributed sensing—where communication range is a controllable design parameter.
Despite these advances, a mechanistic understanding of how information-sharing range determines group-level outcomes remains lacking. The central goal of this study is to understand how social learning range shapes collective foraging. Specifically, we address the following research questions (RQs):
RQ1. How does the social learning parameter, $\rho$, mediate the optimal resolution of the tradeoff between exploration and exploitation in collective search, and how does the optimum $\rho$ vary with factors such as target distribution and group size?
RQ2. How does $\rho$ influence other collective characteristics such as temporal stability and equity of resources found among individuals, as well as their network connectivity?
Here, we develop a collective foraging model that integrates ARS with range-limited social learning. Individuals switch among three behavioral modes—exploration, exploitation, and targeted movement—regulated by a single communication-distance parameter $\rho$, which sets how far social signals propagate. This formulation allows us to quantify how $\rho$ jointly shapes (i) mean search efficiency, (ii) the temporal stability and equity of resource intake, (iii) network connectivity among agents, and the emergent interaction structure underlying these outcomes. We then extend the same framework to risky environments by introducing negative targets that generate avoidance cues shared through the same social-learning rule.
Our results show that $\rho$ induces a non-monotonic performance landscape: small $\rho$ preserves independent exploration but limits coordination, intermediate $\rho$ balances exploration and exploitation to maximize efficiency, and large $\rho$ promotes homogeneous behavior that increases equity but reduces efficiency and produces bursty intake dynamics. Mechanistically, this optimum coincides with an intermediate level of emergent connectivity in the proximity network induced by social learning. Additionally, when negative targets are introduced, the optimal $\rho$ shifts toward larger values as avoidance becomes more valuable; moreover, when penalties are ignored, randomly distributed negative cues can increase efficiency by pruning unproductive search regions. Together, these findings provide a minimal mechanistic framework for how information-sharing range shapes collective foraging, with implications for biological groups and engineered swarms.
@@
Figure 1: A schematic of the collective search model. Agents begin by performing independent Lèvy walks with $\mu=1.1$ (exploration, shown in purple) until they detect a target. Upon detection, the agent ceases social learning and switches to a Lèvy walk with $\mu=3$ (exploitation, shown in blue) within a circular region of radius $R$ centered at the detected target, which is updated over time. Meanwhile, a social signal emitted by this agent attracts other agents within a region of radius $\rho$ to perform targeted walk towards the detected target (shown in orange). Agents outside this region do not receive the social signal and maintain their search mode.
@@ Results
We organize the Results around the two research questions posed in the Introduction. We first address RQ1 by quantifying how the social-learning range $\rho$ mediates the exploration–exploitation trade-off in collective search and produces an intermediate $\rho$ that maximizes collective efficiency. We then address RQ2 by characterizing how $\rho$ reshapes collective properties beyond mean efficiency—including temporal stability of target intake, equity of individual rewards, and the emergent proximity-network connectivity that underlies information flow and coordination. Finally, building on the mechanisms identified in RQ1 and RQ2, we present a post-hoc application in which the same framework is extended to a risky environment (containing negative targets), illustrating how the role of $\rho$ generalizes under additional task constraints. To this end, we simulate the group foraging model on a periodic domain while sweeping $\rho$ from 0 to $0.7$ (the maximum feasible range in our geometry). We vary the number of patches while holding the total number of positive targets fixed at $6000$. To model risky environments we introduce negative targets that reduce efficiency; upon encountering a negative target, an agent broadcasts its location within radius $\rho$, and informed agents exhibit active avoidance behavior. Further details of the simulation setup are provided in Methods.
@@
The group efficiency, defined as $\eta=N_{collected}/N_{positive}T$, quantifies the rate at which agents collect positive targets. Here, $N_{collected}$ and $N_{positive}$ are the number of targets collected by agents and the total number of targets within the domain, respectively. $T$ is the total search time. As the social learning range ($\rho$) increases, it first increases and then decreases, with an optimal $\rho$ that yields the highest efficiency. (Fig. 2A). This phenomenon is robust across different numbers of patches, whether we fix the total number of targets or the number of targets per patch (see Supplementary Fig. 4). It can also be observed for target distributions with $\beta=1.1$ (see Supplementary Fig. 2) and $\beta=3$ (see Supplementary Fig. 3), as well as different group sizes when $N=10$ (see Supplementary Fig. 5). This non-monotonic dependence is consistent with previously identified trade-offs between exploration and exploitation and between individual search and social learning [garg2022individual]. Here, our explicit definition of exploration and exploitation provides a clearer mechanistic account of the emergence of an optimal $\rho$.
@@
Agents in our model exhibit three modes: Lèvy walk with $\mu=1.1$, Lèvy walk with $\mu=3$, and targeted walks triggered by social learning. We interpret $\mu=1.1$ mode as global exploration, $\mu=3$ mode as area restricted exploitation, and the targeted walk as the operational cost associated with exploitation. Agents can adaptively switch between the three modes as needed. We measure the time spent on each mode, as shown in Fig. 2B. When $\rho=0$ (no social learning), targeted walks do not occur, and agents search independently, spending approximately 90% of their time on exploration. After encountering a target, an agent shifts from exploration to exploitation, which contributes to the remaining 10% of time spent on exploitation. As $\rho$ increases, targeted walks become more prevalent, triggering greater exploitation. Consequently, the time ratio of exploration decreases as $\rho$ rises. At the optimal $\rho$, the time ratio of exploitation reaches its maximum. This reflects an elegant balance between exploration and exploitation: enough exploitation sustains high current efficiency, while just enough exploration continually discovers new patches, preserving future opportunities and, in turn, overall efficiency. As $\rho$ increases further, targeted walks become more prevalent. However, the time ratio of exploitation decreases, indicating that an excessively high proportion of targeted walks leads to redundant search effort and reduces the exploitation ratio. This occurs because the number of targets is limited in each patch. When social signals attract most of the agents to a detected target, early-arriving agents can still collect targets, whereas late-arriving agents who travel for extended periods can only collect a few targets. Additionally, the remaining targets may be insufficient to support sustainable exploitation, causing late-arriving agents to revert to exploration soon. In other words, the social signal breaks the exploration of those late-arriving agents (far from the social signal) and lets them restart exploration from a region where most of the targets are consumed. Thus, an excessively large $\rho$, causes a high proportion of targeted walks, raises the cost and reduces the benefits of exploitation, disrupts the exploration process, and forces agents to explore regions where most targets are already consumed, thereby diminishing efficiency.
@@
To investigate the role of social learning and the resulting targeted walk, we construct a phase diagram in 2C, plotting the counts of targeted walks normalized to the total time in each case, against the averaged length of targeted walks. In this diagram, a higher count of targeted walks signifies increased interaction with social signals, which promotes exploitation and enhances efficiency. Conversely, a larger average length of targeted walks represents high cost, as it is the distance agents must travel before they can begin detecting a resource, which reduces overall efficiency. The slope of the curves in this phase diagram, therefore, represents the ratio of benefits to costs, specifically the gain from increased exploitation per unit of cost. Based on this, we hypothesize that the point where the slope of each curve reaches its maximum corresponds to the most efficient scenario. At this point, the benefit gained from targeted walks increases most rapidly relative to their cost. To verify this hypothesis, we compute the derivative of the curves from the phase diagram as shown in Fig. 2D, which reveals the slope. From the figure, a maximum value of the derivative represents the most efficient scenario. This derivative is smoothed using a Gaussian filter with $\sigma=1$ to reduce the noise. Together, these results answer RQ1 by demonstrating that performance is non-monotonic in $\rho$: intermediate-range social learning maximizes efficiency by improving exploitation without fully suppressing independent discovery.
Figure 3: Collective search characteristics induced by $\rho$: burstiness, equity, and network connectivity. (A) Time variation of instantaneous efficiency $\eta_{i}$ with $\rho=0.01$, $\rho=0.15$, and $\rho=0.7$ (showing one case). $\eta_{i}$ represents the number of collected targets per time step. (B) Probability density function (PDF) of $\eta_{i}$ indicating that low $\rho\leq 0.2$ limits $\eta_{i}$ due to less social learning, while higher $\rho$ ($0.2<\rho<0.5$) allows $\eta_{i}$ to reach greater values with an exponential distribution. For $\rho\geq 0.5$, the proportion of medium $\eta_{i}$ values decreases, though $\eta_{i}$ can achieve higher values with a heavy-tail distribution. (C) PDF of inter-event time, defining an event as when the instantaneous efficiency $\eta_{i}\geq 1$. The inset figure illustrates the variation of the burstiness parameter $B$ with $\rho$. The optimal $\rho$ corresponds to the lowest B, indicating a steady intake of resources and, consequently, a minimal starvation time. (D) PDF of number of collected targets by each agent, with an inset indicating that the standard deviation (representing equity) decreases as $\rho$ increases. (E) PDF of the average component size in each case $S$. The PDF of $S$ shifts to the right as $\rho$ increases, indicating that the case-averaged component size $\left\langle S\right\rangle$ increases with $\rho$, as shown in the inset. (F) PDF of average number of components in each case $N_{C}$. The PDF of $N_{C}$ shifts to the left as $\rho$ increases, indicating the case-averaged number of components $\left\langle N_{C}\right\rangle$ decreases with increasing $\rho$, as shown in the inset. All data is based on a scenario with 50 agents and $N_{+}=10$.
@@
To highlight the burstiness of $\eta_{i}$, we plot its probability density functions (PDFs), as shown in Fig. 3B. The PDF reveals that a larger $\rho$ increases the proportion of higher $\eta_{i}$ values, resulting from social learning and agent aggregation. To further reveal the bursty pattern existing in the timeseries of instantaneous efficiency, we use the burstiness parameter introduced by Goh and Barabási [goh2008burstiness]. Here, we defined an event as occurring when $\eta_{i}$ exceeds 0, while the interevent time $\tau$ represents the starvation time during which agents do not collect any targets. The burstiness parameter is quantified by $B\equiv(\sigma_{\tau}-m_{\tau})/(\sigma_{\tau}+m_{\tau})$, where $m_{\tau}$ and $\tau_{m}$ are the mean and standard deviation of $\tau$, respectively. Fig. 3C shows the PDFs of interevent times, where cases with high and low $\rho$ exhibit a larger proportion of starvation time, reflecting greater burstiness and leading to lower efficiency. The inset in Fig. 3C plots the burstiness parameter $B$ as a function of $\rho$, where $\rho$ values with lower $B$ correspond to the optimal $\rho$. Thus, the analysis of the time series and the derived burstiness provides another perspective on the optimal $\rho$, which not only exhibits the highest efficiency but also steady intake of resources and reduced burstiness in starvation time.